Consider the following (illustrative) life table

Age lx

0. 100

  1. 85
  2. 75
  3. 50
  4. 45
  5. 30
  6. 57
  7. 0

Calculate the median lifetime and the

expected remaining lifetime of a life aged 2 for this population.
Answer #1

The lifetimes sorted in ascending order are:

0

30

45

50

57

75

85

Total number of observations(n)=7

Since n is odd, median is (n+1)/2th observation

= (7+1)/2th observation

=8/2th observation

=4th observation

=50 years

So median lifetime is 50 years.

Average lifetime=(0+30+45+50+57+75+85)/7

=342/7

=48.85 which can be rounded to 49

So expected remaining lifetime of a life aged 2 for this population is 49-2=47 years